# CRAFT: Calibrated Reasoning with Answer-Faithful Traces via Reinforcement Learning for Multi-Hop Question Answering

### v1

@@ Abstract
Retrieval-augmented generation (RAG) is widely used to ground Large Language Models (LLMs) for multi-hop question answering. Recent work mainly focused on improving answer accuracy via fine-tuning and structured or reinforcement-based optimization. However, reliable reasoning in response generation faces three challenges: 1) Reasoning Collapse. Reasoning in multi-hop QA is inherently complex due to multi-hop composition and is further destabilized by noisy retrieval. 2) Reasoning–answer inconsistency. Due to the intrinsic uncertainty of LLM generation and exposure to evidence–distractor mixtures, models may produce correct answers that are not faithfully supported by their intermediate reasoning or evidence. 3) Loss of format control. Traditional chain-of-thought generation often deviates from required structured output formats, leading to incomplete or malformed structured content. To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a Group Relative Policy Optimization (GRPO) based reinforcement learning framework that trains models to perform faithful reasoning during response generation. CRAFT employs dual reward mechanisms to optimize multi-hop reasoning: deterministic rewards ensure structural correctness while judge-based rewards verify semantic faithfulness. This optimization framework supports controllable trace variants that enable systematic analysis of how structure and scale affect reasoning performance and faithfulness. Experiments on three multi-hop QA benchmarks show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales, with the CRAFT model achieving competitive performance with closed-source LLMs across multiple reasoning trace settings.
@@ 1 Introduction
Retrieval-augmented generation (RAG) is widely used for multi-hop question answering (QA), where large language models (LLMs) generate answers grounded in retrieved candidate documents. A typical multi-hop QA task requires synthesizing information across multiple pieces of evidence that may be temporally, conceptually, or causally linked, while filtering out irrelevant or misleading distractors introduced by retrieval, calling for explicit reasoning traces that faithfully connect the supporting evidence to the final answer. As illustrated in Figure 1, reliable response generation in RAG systems faces three intertwined difficulties. First, multi-hop reasoning is inherently difficult, as it requires composing information across multiple dependent steps, and this difficulty is further amplified by noisy retrieval that introduces misleading or irrelevant documents. Second, generation by large language models is intrinsically uncertain, and under mixed evidence–distractor contexts, models may additionally exploit shortcut behaviors, producing answers that are weakly supported by the underlying reasoning or evidence. Third, models often fail to reliably follow required output formats during generation, leading to incomplete or malformed structured outputs and limiting controllability of the produced reasoning traces. These issues complicate the learning and assessment of robust reasoning in RAG systems.
@@
Researchers have explored various approaches to improve accuracy in LLM-based multi-hop QA tasks, which can be broadly categorized along three complementary dimensions. (1) Explicit reasoning without guaranteed faithfulness. Early approaches encourage step-by-step reasoning through supervised fine-tuning (SFT), chain-of-thought prompting, and decomposition-based formulations Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Yao et al. (2022). While effective in improving answer accuracy, the resulting reasoning traces are not guaranteed to be grounded in the supporting evidence, especially under noisy retrieval, where models may bypass relevant documents and exploit spurious shortcuts. Trivedi et al. (2020); Press et al. (2023); Lanham et al. (2023); Turpin et al. (2023). (2) Post-hoc grounding without structured reasoning. Another line of work promotes evidence grounding by introducing citations, evidence constraints, or verification signals Gao et al. (2023); Wei et al. (2025); Li et al. (2025a); Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025). But these approaches often apply grounding after generation or only at the answer level, leaving intermediate reasoning traces free-form and unstructured. (3) Optimization of LLMs with coarse-grained rewards. More recent approaches explore reinforcement learning to optimize reasoning and answer generation in retrieval-augmented QA systems Asai et al. (2024); Li et al. (2025b); Chen et al. (2025). These methods typically rely on coarse outcome-level rewards tied to final answer correctness, providing limited supervision over intermediate reasoning and faithfulness properties, and allowing correct answers to be produced without aligned or faithful reasoning.
To address these challenges, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) for the response generation stage of RAG-based LLMs in multi-hop QA under noisy retrieved candidates. CRAFT trains generators to produce structured, machine-auditable reasoning traces, enabling systematic verification of reasoning faithfulness to retrieved evidence and generated answers. Our study consists of two aspects: 1) RL training with fine-grained reward supervision. Given a question and a retrieved candidate set containing both evidence and distractors, CRAFT optimizes reasoning trace generation using group-based reinforcement learning. Training is guided by a decomposed reward function that combines (i) deterministic rewards enforcing machine-checkable constraints, such as output structure validity, citation-set correctness when applicable, and answer matching, with (ii) judge rewards from a high-capacity LLM-as-a-judge that assess semantic faithfulness, including consistency across reasoning stages (e.g., planreasonanswer) and grounding to retrieved evidence. This design provides targeted supervision over unlabeled but critical semantic properties of multi-hop reasoning. 2) Controllable trace space and scale study. CRAFT defines a controllable space of reasoning trace variants by composing trace components (<plan>, <gold_docs>, <reason>, <answer>) along an auditability spectrum. This design enables systematic analysis of how trace structure influences performance and faithfulness, as well as how model scale interacts with trace complexity and audit constraints. Our contributions are three-fold:
Dual reward mechanisms for faithful multi-hop reasoning optimization. We propose CRAFT, a GRPO-based RL framework combining deterministic rewards for verifiable metrics and judge rewards for semantic faithfulness.
Systematic analysis of controllable reasoning traces. We study how different combinations of trace components affect model performance, and characterize how model scale shapes the preferred trace structures and their performance boundaries in small LLMs.
Experimental results demonstrate that CRAFT consistently improves answer accuracy and the faithfulness of generated reasoning traces across model scales. Notably, the CRAFT-trained 7B model achieves competitive performance with closed-source LLMs across different reasoning trace configurations on three multi-hop QA benchmarks.
2 Related Works
2.1 Faithful Reasoning in LLMs
Chain-of-thought (CoT) prompting Wei et al. (2022) improves reasoning in language models, but recent work questions whether these traces faithfully reflect the model’s actual decision process. Lanham et al. (2023) found that CoT faithfulness decreases with model scale, while Turpin et al. (2023) showed that CoT can be influenced by spurious input features; in safety-critical LLM-integrated systems, such unconstrained outputs undermine response reliability Zhang et al. (2025). Recent evaluation methods include Tutek et al. (2025)’s faithfulness measurement via unlearning reasoning steps, Arakelyan et al. (2025)’s FLARE framework using logic programming for verifiable reasoning, and Sui et al. (2025)’s approach anchoring responses to structured knowledge graph paths. However, these methods largely rely on post hoc analysis rather than enforcing machine-checkable structure during training.
2.2 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) Lewis et al. (2020) grounds language models in retrieved documents. For multi-hop reasoning, Trivedi et al. (2023) proposed IRCoT with interleaved retrieval, while recent work includes HopRAG Liu et al. (2025a) using passage graphs, DualRAG Cheng et al. (2025) combining reasoning-augmented querying with knowledge aggregation, and OPERA Liu et al. (2025b) coordinating planning and retrieval via RL. On generation, Gao et al. (2023) trained models for citations, while InstructRAG Wei et al. (2025) and RAG-DDR Li et al. (2025a) optimize generation with reward signals. Despite these advances, most approaches prioritize retrieval quality or answer-level grounding rather than producing structured reasoning traces.
2.3 Reinforcement Learning for LLMs
Reinforcement learning from human feedback (RLHF) Ouyang et al. (2022) aligns language models via policy optimization like PPO Schulman et al. (2017), DPO Rafailov et al. (2023), and RRHF Yuan et al. (2023). GRPO Shao et al. (2024) removes the need for a critic by normalizing rewards within sampled groups, enabling efficient training for reasoning models like DeepSeek-R1 Guo et al. (2025). For RAG, Chen et al. (2025) applied multi-agent RL, while R3-RAG Li et al. (2025b) jointly optimizes reasoning and retrieval with outcome and process rewards. However, most methods use scalar rewards, making it hard to provide targeted supervision over structured reasoning.
@@ 2.4 Multi-hop Question Answering
Multi-hop QA benchmarks include HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). Press et al. (2023) identified a compositionality gap in models, while Trivedi et al. (2020) showed they exploit shortcuts rather than genuine reasoning. Methods like Decomposed Prompting Khot et al. (2023), Least-to-Most Zhou et al. (2023), and ReAct Yao et al. (2022) improve performance. Recent work has focused on understanding reasoning capabilities: Yao et al. (2025) found that models can learn implicit multi-hop reasoning but require exponentially more training data as reasoning depth increases, while Yu et al. (2025) proposed back attention to enhance latent multi-hop reasoning. For evaluation, Lee et al. (2025) introduced GRADE, which generates difficulty-controlled QA pairs based on reasoning depth and retrieval difficulty. Despite these advances, most approaches lack machine-checkable grounding to verify that reasoning steps are supported by evidence.
@@ 3.1 Problem Formulation
We study the response generation stage of RAG-based LLMs for multi-hop QA tasks, where each instance consists of a question and a retrieved candidate document set that mixes supporting evidence with distractors. Answering typically requires reasoning over multiple supporting documents indexed by , while distractors introduce spurious shortcuts. Our goal is to train models that generate not only correct answers but also auditable reasoning traces—structured outputs in XML-style format (<plan>, <gold_docs>, <reason>, <answer>) that expose the derivation process. Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding. This formulation transforms the learning problem from maximizing a sparse answer reward to optimizing a composite objective over the structure, process, and outcome of reasoning.
@@ 3.2 Overview
As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) a CRAFT Interface that defines structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
CRAFT Interface.
We define a family of XML-based output templates spanning an auditability spectrum. As detailed in Algorithm 1 (upper part), the full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning. Intermediate variants (CRAFT–CRAFT) selectively ablate <plan> or <gold_docs>, while CRAFT (Answer-Only) serves as a minimal baseline. This design enforces a chain of faithfulness: —any hallucination breaks at least one link and is detectable.
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Open-source | ||||||||||
| Qwen2.5-0.5B Team (2024) | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
| Qwen2.5-1.5B Team (2024) | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
| Qwen2.5-3B Team (2024) | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
| Qwen2.5-7B Team (2024) | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
| Qwen3-4B Team (2025) | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
| Qwen3-30B Team (2025) | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
| API | ||||||||||
| GPT-5-mini OpenAI (2025) | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
| Deepseek-V3.2 DeepSeek-AI (2025) | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
| Gemini-2.5-Flash Google DeepMind (2025) | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
| SFT | ||||||||||
| Qwen2.5-0.5B Team (2024) | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
| Qwen2.5-1.5B Team (2024) | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
| Qwen2.5-3B Team (2024) | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
| Qwen2.5-7B Team (2024) | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
| Ours (RL) | ||||||||||
@@ 3.3 CRAFT Training
where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing . The key design choice is the total reward , which we decompose into four components:
@@
We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1 for the detailed procedure). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
@@
where and depends on the trace variant (e.g., plan/citation checks apply only when the corresponding fields exist). Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in .
Why no direct supervision for Plan/Reason text?
Unlike evidence selection (discrete indices) or final answers (short spans), the intermediate <plan> and <reason> are open-ended natural language generations with high semantic variance. Direct supervision (e.g., via SFT on teacher traces) often leads to brittle imitation or format collapse on out-of-distribution data. Instead of enforcing a rigid “gold plan,” our rewards focus on functional correctness via consistency checks (): a plan is valid if it is followed; a reason is valid if it is grounded. This allows the model to discover its own optimal reasoning paths within the auditable structure.
4 Experiment
@@
| EM | F1 | Faithfulness | EM | F1 | Faithfulness | EM | F1 | Faithfulness | ||
| Qwen2.5-7B | ||||||||||

### v2

# CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering

@@ Abstract
Retrieval-augmented large language models, when optimized with outcome-level rewards, can achieve strong answer accuracy on multi-hop questions. However, under noisy retrieval, models frequently suffer from right-answer-wrong-reason failures: they may exploit spurious shortcuts or produce reasoning traces weakly grounded in the supporting evidence. Furthermore, the lack of structured output control prevents reliable auditing of the underlying reasoning quality. To address this, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework for the response generation stage of retrieval-augmented multi-hop question answering. CRAFT trains models to produce structured reasoning traces with configurable levels of auditability (e.g., by selectively retaining planning, evidence citation, or reasoning steps). Training combines two complementary forms of supervision: deterministic rewards enforce verifiable constraints, including format compliance, answer correctness, and citation-set validity, while a judge-based reward audits semantic faithfulness by evaluating reasoning consistency and evidence grounding. Experiments show that CRAFT improves both answer accuracy and reasoning faithfulness across model scales. Notably, semantic judge-based rewards improve answer accuracy rather than compromise it, enabling CRAFT to achieve performance competitive with strong closed-source models.
CRAFT: Calibrated Reasoning with Answer-Faithful Traces for Multi-Hop Question Answering
Yu Liu1, 2††thanks: Equal contribution., Wenxiao Zhang311footnotemark: 1, Diandian Guo1, 2, Cong Cao1††thanks: Corresponding authors., Fangfang Yuan1, Qiang Sun3, Yanbing Liu1, 2, Jin B. Hong3, Zhiyuan Ma422footnotemark: 2 1Institute of Information Engineering, CAS, 2University of Chinese Academy of Sciences, 3The University of Western Australia, 4Huazhong University of Science and Technology caocong@iie.ac.cn, mzyth@hust.edu.cn
@@ 1 Introduction
Retrieval-augmented generation (RAG) grounds large language models (LLMs) in external documents and is widely used for multi-hop question answering (QA), where answers require composing information across multiple passages Lewis et al. (2020). However, a correct answer does not imply faithful reasoning: models can exploit spurious shortcuts—citing irrelevant documents, confusing entities, or following logically inconsistent chains—yet still reach the gold answer (Figure 1, top). Such right-answer-wrong-reason failures are invisible to standard metrics that evaluate only the final output. The problem is especially acute in multi-hop QA, which demands compositional inference over dependent steps (e.g., resolving a bridge entity and linking its attributes to downstream facts) and is further destabilized by distractors that share surface overlap with the query Trivedi et al. (2020); Press et al. (2023).
@@
Prior work addresses aspects of this problem from three directions, but each leaves significant gaps. (1) Explicit reasoning without guaranteed faithfulness. Chain-of-thought and decomposition-based methods encourage step-by-step reasoning Wei et al. (2022); Khot et al. (2023); Zhou et al. (2023); Yao et al. (2022), but the resulting traces are not grounded in the retrieved evidence; under noisy retrieval, models may bypass relevant documents and produce plausible-looking but unfaithful rationales Lanham et al. (2023); Turpin et al. (2023). (2) Post-hoc grounding without structured reasoning. Citation and verification mechanisms introduce evidence signals such as document-level citations or consistency checks Gao et al. (2023); Wei et al. (2025); Li et al. (2025a); Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025), yet these are typically applied after generation or enforced only at the answer level, leaving intermediate reasoning steps unstructured and difficult to audit at scale. (3) Outcome-level optimization without process supervision. Reinforcement learning approaches optimize reasoning and generation Asai et al. (2024); Li et al. (2025b); Chen et al. (2025), but most rely on coarse outcome-level rewards tied solely to final answer correctness, providing weak credit assignment for multi-hop traces and no supervision over intermediate reasoning faithfulness.
To address these limitations, we propose CRAFT (Calibrated Reasoning with Answer-Faithful Traces), a reinforcement learning framework based on Group Relative Policy Optimization (GRPO) Shao et al. (2024) for the response generation stage of retrieval-augmented multi-hop QA. CRAFT defines a family of structured output templates that decompose generation into auditable stages (planning, evidence selection, step-by-step reasoning, and answer extraction), forming trace variants with configurable levels of auditability (Figure 1, bottom). Within each variant, these stages form a chain of faithfulness where each component constrains the next, so that unfaithful reasoning is more likely to violate at least one verifiable link. Training combines two complementary forms of supervision: deterministic rewards enforce format validity, citation-set correctness, and answer matching, while a judge-based reward from a high-capacity LLM audits semantic faithfulness by verifying cross-stage consistency (planreason, citationreason, reasonanswer) and grounding individual claims against the retrieved evidence. This extends process-level supervision beyond rule-based objectives to open-ended semantic properties (e.g., reasoning consistency and claim verifiability). The judge operates on binary decomposed criteria to limit bias propagation, and its reliability is validated against human annotations (§4.5). Together, these designs enable the model to learn not just what to answer but how to reason faithfully. Our contributions are as follows:
We propose CRAFT, an RL framework for retrieval-augmented QA that generates controllable trace variants (CRAFT–CRAFT), improving performance and reasoning auditability.
We introduce a dual-reward design that combines deterministic rewards for verifiable constraints with judge-based rewards for reasoning faithfulness, extending process-level supervision from rule-based objectives to open-ended semantic properties.
Experiments across multiple model scales show that CRAFT substantially improves reasoning faithfulness after training, and further reveal that semantic judge-based rewards can improve answer accuracy rather than compromise it. CRAFT achieves faithfulness and accuracy comparable to larger closed-source models.
2 Related Work
2.1 Faithful Reasoning and Grounding in RAG
CoT prompting Wei et al. (2022) improves reasoning but may not faithfully reflect the model’s decision process Lanham et al. (2023); Turpin et al. (2023); Zhang et al. (2025). Post-hoc faithfulness methods Tutek et al. (2025); Arakelyan et al. (2025); Sui et al. (2025) do not enforce structure during training. RAG Lewis et al. (2020) grounds models in documents, with multi-hop extensions including IRCoT Trivedi et al. (2023), HopRAG Liu et al. (2025a), and OPERA Liu et al. (2025b). Generation grounding via citations Gao et al. (2023) or reward signals Wei et al. (2025); Li et al. (2025a) primarily targets retrieval quality or answer-level grounding rather than structured reasoning with training-time faithfulness enforcement.
2.2 Reinforcement Learning for Reasoning
RLHF Ouyang et al. (2022) aligns models via PPO Schulman et al. (2017), DPO Rafailov et al. (2023), and RRHF Yuan et al. (2023); GRPO Shao et al. (2024) removes the critic via group-wise normalization for reasoning models Guo et al. (2025). Process rewards improve exploration efficiency Setlur et al. (2025) but require domain-specific verifier data. For RAG, multi-agent RL Chen et al. (2025) and joint reasoning-retrieval optimization Li et al. (2025b) have been explored, yet most methods use scalar answer-correctness rewards. LLM-as-a-Judge Zheng et al. (2023) offers scalable reward signals but introduces biases Chen et al. (2024); Gu et al. (2024); mitigations include binary criteria, rubric-based scoring, and human validation Adlakha et al. (2024); Bavaresco et al. (2024). CRAFT adopts these strategies with binary pass/fail judgments across decomposed dimensions, complemented by deterministic rewards.
@@ 2.3 Multi-hop Question Answering
Multi-hop QA benchmarks Yang et al. (2018); Ho et al. (2020); Trivedi et al. (2022) require composing information across multiple documents, yet models exhibit a compositionality gap Press et al. (2023) and exploit shortcuts Trivedi et al. (2020). Decomposition-based Khot et al. (2023); Zhou et al. (2023) and agentic Yao et al. (2022) methods improve performance, while recent work examines implicit reasoning Yao et al. (2025), back attention Yu et al. (2025), and difficulty-controlled evaluation Lee et al. (2025). Most approaches lack machine-checkable grounding to verify reasoning faithfulness.
@@ 3.1 Problem Formulation
We study the response generation stage of RAG-based LLMs for multi-hop QA tasks, where each instance consists of a question and a retrieved candidate document set that mixes supporting evidence with distractors. Answering typically requires reasoning over multiple supporting documents indexed by , while distractors introduce spurious shortcuts. Our goal is to train models that generate not only correct answers but also auditable reasoning traces — structured outputs in XML-style format (<plan>, <gold_docs>, <reason>, <answer>) that expose the derivation process. Formally, given and a gold answer , the model outputs a trace containing components (plan), (citation set), (reasoning), and (answer), satisfying: (i) format parsability , (ii) answer correctness , and (iii) trace faithfulness , where enforces internal consistency and evidence grounding.
@@ 3.2 Overview
As illustrated in Figure 2, CRAFT consists of three tightly integrated components: (1) CRAFT Trace Variants that define structured output templates (top), (2) a Decomposed Reward Auditor that provides multi-signal supervision (center), and (3) a Faithfulness Audit Principle that enables machine-checkable reasoning verification (right).
Trace Templates.
We define a family of XML-based output templates spanning an auditability spectrum. The full template (CRAFT) decomposes model output into four fields forming an auditable chain: <plan> declares the reasoning blueprint (sub-question decomposition); <gold_docs> commits to an evidence boundary before reasoning begins; <reason> must cite only within this boundary while addressing the plan; <answer> must be logically entailed by the reasoning. Intermediate variants selectively ablate components: CRAFT removes <plan>, CRAFT removes <gold_docs>, CRAFT retains only <reason>+<answer>, and CRAFT (Answer-Only) serves as a minimal baseline. This design enforces a chain of faithfulness: — any hallucination breaks at least one link and is detectable. The trace template also determines which rewards are applicable: variants containing <gold_docs> activate , while variants with reasoning fields activate .
@@ 3.3 CRAFT Training
where is the standard deviation of and is a small constant for numerical stability; then performs a critic-free policy update by maximizing:
The key design choice is the total reward , which we decompose into four components:
@@ 3.3.2 Judge Reward
We penalize hallucinated reasoning by verifying internal consistency and citation grounding (Algorithm 1). Let denote extracting the trace fields from . We employ a high-capacity LLM judge to compute:
@@
where and depends on the trace variant (e.g., plan/citation checks apply only when the corresponding fields exist). Concretely, the judge checks: (i) planreason (), (ii) citation-setreason (), (iii) reasonanswer (), and (iv) grounding (), i.e., key claims in are supported by the cited document text in . Each check yields a binary score; is their average. The binary, dimension-decomposed design limits the scope for judge error compared to holistic scoring. The judge model (Qwen3-Max) is distinct from the trained policy to avoid self-enhancement bias. Its reliability is validated against human annotations (§4.5), and cross-judge evaluation with an independent model confirms that results are not specific to the choice of judge (Appendix C).
| Method | Type | MuSiQue | HotpotQA | 2WikiMHQA | ||||||
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
| Qwen2.5-0.5B | Pre-Trained | 0.00 | 0.00 | 0.25 | 1.55 | 5.74 | 1.70 | 6.00 | 9.93 | 1.55 |
| Qwen2.5-1.5B | Pre-Trained | 0.80 | 3.73 | 5.40 | 9.80 | 17.41 | 14.70 | 10.40 | 16.08 | 9.10 |
| Qwen2.5-3B | Pre-Trained | 12.40 | 19.50 | 24.35 | 33.00 | 44.29 | 57.05 | 30.55 | 38.80 | 45.25 |
| Qwen2.5-7B | Pre-Trained | 36.85 | 47.79 | 58.25 | 56.95 | 71.55 | 83.65 | 57.70 | 66.72 | 76.85 |
| Qwen3-4B | Pre-Trained | 23.75 | 32.32 | 52.35 | 51.30 | 65.64 | 87.50 | 53.85 | 64.16 | 86.60 |
| Qwen3-30B | Pre-Trained | 40.00 | 49.03 | 68.35 | 59.50 | 74.25 | 92.75 | 56.30 | 66.24 | 90.35 |
[…]
| GPT-5-mini | API | 53.45 | 68.39 | 83.65 | 61.30 | 80.09 | 96.35 | 70.55 | 81.01 | 95.95 |
| Deepseek-V3.2 | API | 53.00 | 67.67 | 77.24 | 62.85 | 79.52 | 94.55 | 68.70 | 77.67 | 96.50 |
| Gemini-2.5-Flash | API | 52.25 | 63.55 | 77.20 | 63.10 | 78.84 | 93.25 | 70.15 | 79.18 | 94.30 |
[…]
| Qwen2.5-0.5B | Trained | 0.00 | 0.00 | 0.11 | 0.81 | 2.87 | 0.77 | 2.85 | 4.72 | 0.68 |
| Qwen2.5-1.5B | Trained | 7.95 | 14.64 | 10.90 | 34.55 | 45.96 | 33.05 | 34.80 | 41.97 | 34.65 |
| Qwen2.5-3B | Trained | 19.45 | 26.71 | 27.90 | 49.00 | 63.24 | 67.60 | 49.50 | 58.12 | 66.45 |
| Qwen2.5-7B | Trained | 32.35 | 43.38 | 50.85 | 57.75 | 72.04 | 83.50 | 62.65 | 71.32 | 78.95 |
[…]
@@
| EM | F1 | Faith. | EM | F1 | Faith. | EM | F1 | Faith. | ||
[…]
@@
Why no direct supervision for Plan/Reason?
Unlike (discrete indices) and (short spans), and are open-ended generations with high semantic variance, making direct supervision via teacher traces brittle and prone to format collapse. Instead, we supervise functional correctness: is valid iff follows it (); is valid iff it is grounded (). This lets the model discover optimal reasoning paths within the auditable structure without imitating a fixed trace.
4 Experiments